1/1

Chānglíng Wáng ,  Chānglíng Wáng ,
Cǎi lián qǔ  The green Skirt of the Girl’s

tradukita de Shirley Yiping Zhang

Héyè luóqún yīsè cái fúróng xiàngliǎn liǎnbiā kāi Luàrù chízhōng kànbújiàn, wéngē shǐjué yǒurēnlái.

The green skirts of the girls’ seem as being cut by the lotus leaves, The pretty lotuses are blossoming to the girls’ pink faces...

Both of them mix into the lotus pool and the girls could not been seen at all... We just realize that some poeple are coming after hearing their beautiful songs...

Verkinto de tiu ĉi ĉina poemo estas Chānglíng Wáng .

Arg-908-1811 (2013-07-11 23:39:45)

Tiu ĉi poemo troviĝas en http: // www. ebridge. cn/ new/ music/ shirley1. php? sno= 531 . Pri la poeto vidu la retejon http: // en. wikipedia. org/ wiki/ Wang_Changling .

Traduko de la ĉina poemo “Cǎi lián qǔ” de Chānglíng Wáng en la Anglan de Shirley Yiping Zhang.

Arg-908-1812 (2013-07-11 22:17:14)

http://www.poezio.net/version?poem-id=908&version-id=1811,1812